Mikaela1509:

Explain how to solve the following equation. 1.5x = 6 How would u explain this?

carmelle:

To find the value of x, you simply have to isolate the variable (basically meaning getting x by itself). To do so, you just divide both sides by 1.5 1.5 ÷ 1.5 = ? 6 ÷ 1.5 = ?

Mikaela1509:

Ohhhh! thats cool!

carmelle:

do you understand?

Mikaela1509:

Yep i do understand

carmelle:

alright, thats great. so what did you get as your final answer?

Israel22:

1.5 ÷ 1.5 6 ÷ 1.5 =